# Real-Time Object Detection Meets DINOv3

### v2

@@ Abstract
Benefiting from the simplicity and effectiveness of Dense O2O and MAL, DEIM has become the mainstream training framework for real-time DETRs, significantly outperforming the YOLO series. In this work, we extend it with DINOv3 features, resulting in DEIMv2. DEIMv2 spans eight model sizes from X to Atto, covering GPU, edge, and mobile deployment. For the X, L, M, and S variants, we adopt DINOv3-pretrained / distilled backbones and introduce a Spatial Tuning Adapter (STA), which efficiently converts DINOv3’s single-scale output into multi-scale features and complements strong semantics with fine-grained details to enhance detection. For ultra-lightweight models (Nano, Pico, Femto, and Atto), we employ HGNetv2 with depth and width pruning to meet strict resource budgets. Together with a simplified decoder and an upgraded Dense O2O, this unified design enables DEIMv2 to achieve a superior performance–cost trade-off across diverse scenarios, establishing new state-of-the-art results. Notably, our largest model, DEIMv2-X, achieves 57.8 AP with only 50.3M parameters, surpassing prior X-scale models that require over 60M parameters for just 56.5 AP. On the compact side, DEIMv2-S is the first sub-10M model (9.71M) to exceed the 50 AP milestone on COCO, reaching 50.9 AP. Even the ultra-lightweight DEIMv2-Pico, with just 1.5M parameters, delivers 38.5 AP—matching YOLOv10-Nano (2.3M) with $\sim$50% fewer parameters.
@@ 1 Introduction
Among mainstream real-time object detectors, DETR-based approaches have attracted increasing attention for their end-to-end design and the high capacity enabled by Transformers, achieving a more favorable trade-off. Within this paradigm, DEIM has emerged as a strong training framework that has advanced real-time DETRs and delivered leading models in the field. Meanwhile, DINOv3 [21] has demonstrated strong feature representation capabilities across a variety of vision tasks. However, their potential for real-time object detection has not been fully explored.
@@ 2 Method
Our training strategy follows DEIM [7], a fundamental framework designed for fast convergence and high performance. The overall optimization objective is a weighted sum of five components: Matchability-Aware Loss (MAL) [7], Fine-Grained Localization (FGL) Loss [17], Decoupled Distillation Focal (DDF) Loss [17], BBox Loss (L1), and GIoU Loss [19]. The total loss is defined as:

### v3

@@ Abstract
Driven by the simple and effective Dense O2O, DEIM demonstrates faster convergence and enhanced performance. In this work, we extend it with DINOv3 features, resulting in DEIMv2. DEIMv2 spans eight model sizes from X to Atto, covering GPU, edge, and mobile deployment. For the X, L, M, and S variants, we adopt DINOv3-pretrained / distilled backbones and introduce a Spatial Tuning Adapter (STA), which efficiently converts DINOv3’s single-scale output into multi-scale features and complements strong semantics with fine-grained details to enhance detection. For ultra-lightweight models (Nano, Pico, Femto, and Atto), we employ HGNetv2 with depth and width pruning to meet strict resource budgets. Together with a simplified decoder and an upgraded Dense O2O, this unified design enables DEIMv2 to achieve a superior performance–cost trade-off across diverse scenarios, establishing new state-of-the-art results. Notably, our largest model, DEIMv2-X, achieves 57.8 AP with only 50.3M parameters, surpassing prior X-scale models that require over 60M parameters for just 56.5 AP. On the compact side, DEIMv2-S is the first sub-10M model (9.71M) to exceed the 50 AP milestone on COCO, reaching 50.9 AP. Even the ultra-lightweight DEIMv2-Pico, with just 1.5M parameters, delivers 38.5 AP—matching YOLOv10-Nano (2.3M) with $\sim$50% fewer parameters.
@@ 1 Introduction
Among prevailing real-time detectors, DETR-based methods are increasingly favored for their end-to-end nature. Despite the robust feature representations offered by DINOv3 [21] in diverse scenarios, the feasibility and effectiveness of integrating it into DETR-based models have yet to be fully investigated.
@@ 2 Method
The overall optimization objective is a weighted sum of five components: Matchability-Aware Loss (MAL) [7], Fine-Grained Localization (FGL) Loss [17], Decoupled Distillation Focal (DDF) Loss [17], BBox Loss (L1), and GIoU Loss [19]. The total loss is defined as: